FTMW SPECTROSCOPY OF HSiCl

\underline{MASARU FUKUSHIMA}^{a}^{b} and WOLFGANG JÄGER, Department of Chemistry, University of Alberta, Edmonton, AB, Canada, T6G 2G2.

We have observed HSiCl using the technique of Fourier transform microwave spectroscopy. HSiCl was produced in supersonic free expansions by pulsed discharge of (CH_{3})_{3}SiCl. The lowest frequency rotational transition, 1_{01} - 0_{00}, was measured and the hyperfine structures due to the nuclear spins of ^{35}Cl and ^{37}Cl were analyzed. The obtained rotational constants, \frac{1}{2} (B+C), were compared with those obtained by LIF optical spectroscopy^{c}. Using the nuclear electric quadrupole and nuclear spin-rotation coupling constants obtained, the electronic structure of HSiCl was investigated.

aPermanent address : Faculty of Information Sciences, Hiroshima City University, Asa-Minami, Hiroshima 731-3194, Japan.

bSupport by Hiroshima City University President-Designated Grant for Long-Term Research Abroad for Young Researchers is gratefully acknowledged.

cW. W. Haper and D. J. Clouthier, J. Chem. Phys. 106, 9461 (1997).